OpenStudy (anonymous):

This table shows 10 observations of a pair of variables (x,y). The variables x and y are positively correlated, with a correlation coefficient of 0.977. Find the slope, b, of the least squares regression line, y=ax+b, for these data: X 1 2 3 4 5 6 7 8 9 11 Y 2 4 5 8 9 12 13 14 15 16

OpenStudy (anonymous):

.The coefficients "a" and "b" for least squares regression may be found as follows: a = [n∑(xy) - ∑(x)∑(y)]/[n∑(x²) - ∑(x)²] b = [∑(x²)∑(y) - ∑(x)∑(xy)]/[n∑(x²) - ∑(x)²] 1. For the equation y = ax + b: a = [(10)(688) - (56)(98)]/[(10)(406) - (56)²] = 1.5065 b = [(406)(98) - (56)(688)]/[(10)(406) - (56)²] = 1.3636 2. For the equation y = ax + b: a = [(4)(68) - (10)(21)]/[(4)(30) - (10)²] = 3.1000 b = [(30)(21) - (10)(68)]/[(4)(30) - (10)²] = -2.5000

OpenStudy (anonymous):

1. The slope, actually, is a, by the way. y = 1.506a + 1.364 So the slope is 1.506. 2. y = 3.1x - 2.5 The slope is 3.1.